# Marvel: Accelerating Safe Online Reinforcement Learning with Finetuned Offline Policy

### v2

@@ 3 Warm-Start Safe RL with Value Pre-Alignment
As demonstrated in Fig. 1, naively finetuning the offline policy for safe RL would not work well and the finetuned policy shows clear “inertia” in improving its performance: within a long period after online finetuning starts, its cost stays far below the limit, but its reward is quite low and not improving at all. This implies that such a strategy automatically “inherits” the conservatism from offline safe RL and is reluctant to actively explore in order to fully utilize the safe gap below the cost limit. In this section, we delve into the failure of naive finetuning, which points to two unique challenges for policy finetuning in O2O safe RL, i.e., erroneous offline Q-estimations and Lagrange multiplier mismatch. To address these problems, we propose a framework for O2O safe RL, namely warM-stArt safe Reinforcement learning with Value prE-aLignment (Marvel).
@@ 3.1 Pre-Finetune Phase
Challenge I: Erroneous Q-estimations resulted from objective mismatch and offline cost sparsity.  By learning from a fixed dataset without online environment interactions, offline safe RL typically suffers from large extrapolation errors for OOD actions beyond the support of the dataset. A general principle to handle this is to penalize the reward/cost estimations for the OOD actions in such a way that risky explorations outside the dataset are discouraged. Particularly, the optimization of Q-functions in offline safe RL can be captured as follows:
Offline ($\textstyle\text{Offline (}Q\text{):}$: $\min\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[\left(Q(s,a)-\left(r+\gamma\mathbb{E}\left[\max_{a^{\prime}}Q(s^{\prime},a^{\prime})\right]\right)\right)^{2}\right]+\psi\cdot\mathcal{P}(s,a_{OOD}),$
Offline ($\text{Offline (}Q_{c}\text{):}$: $\min\mathbb{E}_{(s,a,c,s^{\prime})\sim\mathcal{D}}\left[\left(Q_{c}(s,a)-\left(c+\gamma\mathbb{E}\left[\max_{a^{\prime}}Q_{c}(s^{\prime},a^{\prime})\right]\right)\right)^{2}\right]-\psi_{c}\cdot\mathcal{P}_{c}(s,a_{OOD}).$
@@
Here $\psi\cdot\mathcal{P}(s,a_{OOD})$ and $\psi_{c}\cdot\mathcal{P}_{c}(s,a_{OOD})$ are the penalty terms. For instance, penalties are introduced in VOCE [18] so as to minimize the expected reward Q-values and maximize the expected cost Q-values for OOD actions. CPQ [19] increases the perceived cost of OOD actions during Q-function and policy updates, while keeping cost below the threshold.
In contrast, the optimization of Q-functions in online safe RL is standard without any penalty terms:
Online ($\textstyle\text{Online (}Q\text{):}$: $\min\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[\left(Q(s,a)-\left(r+\gamma\mathbb{E}\left[\max_{a^{\prime}}Q(s^{\prime},a^{\prime})\right]\right)\right)^{2}\right],$
Online ($\text{Online (}Q_{c}\text{):}$: $\min\mathbb{E}_{(s,a,c,s^{\prime})\sim\mathcal{D}}\left[\left(Q_{c}(s,a)-\left(c+\gamma\mathbb{E}\left[\max_{a^{\prime}}Q_{c}(s^{\prime},a^{\prime})\right]\right)\right)^{2}\right].$
@@
Obviously, Offline and online safe RL have distinct objectives for Q-functions, meaning pretrained Q-functions may not accurately estimate values for state-action pairs encountered during online interactions. As a result, offline policies tend to act overly conservatively, exploring only low-cost regions during online finetuning. However, effective online learning requires identifying state-action pairs with both high rewards and low costs, which necessitates exploring areas with potentially higher costs. This objective mismatch is even more pronounced in O2O safe RL due to the additional Q-function for cost estimation. The sparsity of offline cost leads to a pretrained cost Q-function that predicts low costs for IND state-actions, further limiting exploration during finetuning.
@@
Solution: Value Pre-Alignment.  To address the first challenge, a naive approach is to reevaluate the offline policy in online environments using Monte Carlo simulations, which however introduces additional interaction costs. Motivated by the recent advances in Off-Policy Evaluation (OPE) [28], we borrow the idea from Fitted Q Evaluation [29] to align the offline Q-functions with the online learning objectives for the offline policy, by using the offline dataset before online policy finetuning. In particular, we seek to minimize the following objectives for the reward and cost Q-functions by starting from the pretrained Q-functions from offline learning, respectively:
$\textstyle\mathcal{L}_{Q}^{VPA}(\mu)$ $=\displaystyle\mathbb{E}_{D}\left[\mathcal{L}_{2}\left(Q(s,a;\mu)-(r+\gamma\displaystyle\mathbb{E}_{a^{\prime}\sim\pi_{\theta}(\cdot|s^{\prime})}[\hat{Q}(s^{\prime},a^{\prime};\mu^{\prime})-\alpha^{VPA}\log\pi_{\theta}(a^{\prime}|s^{\prime})])\right)\right]$ (8)
@@
$\mathcal{L}_{Q_{c}}^{VPA}(\mu_{c})$ $=\displaystyle\mathbb{E}_{D}\left[\mathcal{L}_{2}\left(Q_{c}(s,a;\mu_{c})-(c+\gamma\displaystyle\mathbb{E}_{a^{\prime}\sim\pi_{\theta}(\cdot|s^{\prime})}[\hat{Q}_{c}(s^{\prime},a^{\prime};\mu_{c}^{\prime})-\alpha_{c}^{VPA}\log\pi_{\theta}(a^{\prime}|s^{\prime})])\right)\right],$ (9)
where $D$ represents the distribution of $(s,a,s^{\prime})$ in the offline dataset, and $\mathcal{L}_{2}$ denotes the MSE loss. $\mu$ denotes parameters of Q function and $\hat{Q}$ is the target Q function. Entropy terms are introduced in VPA to 1) optimistically estimate rewards to encourage greater exploration during the early stages of finetuning and 2) pessimistically estimate costs to ensure the agent remains cautious about the cost threshold during exploration. Specificity, the entropy terms can result in both higher rewards and costs for state-action pairs with high entropy, where the pretrained policy is ‘uncertain’. We set the coefficient $\alpha$ for the Qc-network lower than $\alpha_{c}$ for the Q-network to encourage higher exploration during the agent’s finetuning process. During the initial stage of online finetuning, the agent prioritizes exploring these high-reward areas, even at a high cost, as reward maximization dominates due to the small Lagrangian coefficient. This process helps refine the Q values through interactions. The offline policy remains unchanged during VPA to preserve the knowledge extracted from offline data. As a result, the agent can potentially explore high cost areas, without being overly conservative.
@@
Table 1: The Spearman’s rank correlation coefficients of the $Q$-value and $Q_{c}$-value are evaluated in the BallCircle and CarRun environments. “Random” refers to rollouts where the offline policy starts from a randomly initialized state-action pair (may not in the offline dataset), while “Dataset” refers to rollouts starting from a state-action pair randomly selected from the offline dataset.
@@
To characterize the performance of VPA in correcting the Q-estimations, we leverage Spearman’s rank correlation coefficient, which measures the strength and direction of a monotonic relationship between two ranked variables. The reason is that the relative ranking of Q-values are more important than the absolute values for policy update. Specifically, given a dataset collected by rolling out the offline policy in the environment, we compare the ranking of learned reward/cost Q-values before and after VPA with that of estimated actual return by using Monte Carlo simulations. A large Spearman’s rank correlation coefficient implies that the distribution of learned Q-values is more aligned with the distribution of true Q-values. As shown in Table 1, it is evident that the coefficient increases significantly after VPA for both reward and cost Q-values, no matter if the offline policy rolls out from a seen state-action pair in the offline dataset or from a randomly selected OOD state-action pair. This clearly demonstrates the effectiveness of VPA in aligning the pretrained Q-functions.
@@ 3.2 Finetune Phase
Challenge II: Lagrange multiplier mismatch.  Conventional value-based online safe RL relies on updating Lagrange multipliers alongside the policy and Q-functions during training, so as to push the overall cost below the limit while striking a right balance between maximizing the reward and minimizing constraint violations. While the policy and Q-functions can benefit from offline pretraining for a warm start, offline safe RL algorithms like CPQ [19] and BEAR-lag [9] cannot accurately estimate Lagrange multipliers with regularizing strengths matching with the cost of the offline policy, e.g., a small Lagrange multiplier is not power enough to push down the policy cost, while a large multiplier prevents active exploration of high-reward state-action pairs. For instance, in the BallCircle environment, the offline pretrained Lagrange multiplier value obtained using the BEAR-lag algorithm is approximately 1500, whereas during online finetuning, the SAC-lag requires a value of only about 0.65. The gap between these values clearly precludes the direct use of offline pretrained Lagrange multipliers. Improper initialization can lead to extensive constraint violations or training stagnation, an issue we term as the Lagrange multiplier mismatch.
@@
On the other hand, as VPA promotes active exploration of high-reward state-actions by optimistically estimating rewards and pessimistically estimating costs, it inevitably increases the risk of exploring high cost state-actions, which in turn amplifies the need for appropriate Lagrange multipliers to quickly reduce the constraint violations. Figure 2 shows the online finetuning performance comparison after VPA in two environments between 1) empirically setting a good initial value for the Lagrange multiplier and 2) setting it to zero, where the traditional dual ascent method is used to update the multiplier. It is clear that a good initial multiplier can manage the cost very well, while the policy with a very small initial multiplier value suffers from large constraint violations and takes a much longer time to reduce the cost below the limit. The results also imply that although VPA aligns the distributions of Q-values, it may introduce high costs for online finetuning, which can be addressed with an appropriate initial Lagrange multiplier.
Solution: Adaptive PID Control.  Clearly, finding a good initial value of the Lagrange multiplier can jointly address the mismatch problem and mitigate the potential risk of VPA. However, achieving this through experimental tuning is challenging, and currently, there is no theory to accurately predict these values. To address this problem, instead of directly finding the best initial value, we take an alternative path by quickly adapting the Lagrange multipliers in an effective manner. Motivated by the recent success of leveraging PID control for updating the multiplier in online safe RL [20, 30, 31], we introduce an adaptive PID control approach specifically tailored for O2O safe RL.
More specifically, compared to standard dual ascent in Eq. 3, PID control [32] offers a different approach to updating the Lagrange multiplier in the following way:
[…]
where $e(t)$ captures the cost violation (errors) at time t, namely the cumulative cost difference from the policy rollout compared to the cost threshold. Here, $K_{p}$ is the proportional gain, corresponding to the instantaneous value of the error; $K_{i}$ is the integral gain to characterize the accumulation of past errors and $K_{d}$ is the differential gain corresponding to the rate of change of the error. In practice, the integral and differential are usually discretized as follows:
[…]
By taking the rate of change into consideration, the PID control can be especially useful in scenarios where costs fluctuate significantly, which however is not sufficient to handle the unique challenges in O2O safe RL. Specifically, to address the objective mismatch and the over conservatism exacerbated by sparse costs in the offline learning, VPA encourages active exploration but increases the risk of high cost. This requires a stronger control strength at the early stage of online finetuning to quickly reduce the cost. As training progresses and the cost is approaching to the limit, a weaker control strength is however preferred to stabilize the learning and keep the cost below the limit without large oscillations. The need of dynamic strength points to the need of adaptive control for O2O safe RL.
Towards this end, we propose an adaptive PID control approach for updating the Lagrange multiplier, which dynamically adjusts the PID control parameters during online finetuning based on the incurred policy cost over a time window of $n$ steps:
@@
$K_{p}\leftarrow\text{clip}\left(K_{p}\cdot\left(1+\alpha\cdot\tanh\left(\frac{\bar{c}-{c_{th}}}{\bar{c}}\right)\right),K_{p_{min}},K_{p_{max}}\right),$ (11)
@@
$K_{i}\leftarrow\text{clip}\left(K_{i}\cdot\left(1+\beta\cdot\frac{\bar{c}-c_{th}}{\bar{c}}\right),K_{i_{min}},K_{i_{max}}\right),~{}K_{d}\leftarrow\text{clip}\left(K_{d}\cdot\left(1+\gamma\frac{\sigma_{c}}{\bar{c}}\right),K_{d_{min}},K_{d_{max}}\right)$ (12)
where the average cost $\bar{c}=\frac{1}{n}\sum_{i=1}^{n}c_{i}$, the standard deviation $\sigma_{c}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(c_{i}-\bar{c})^{2}}$, $\alpha$, $\beta$ and $\gamma$ are hyper-parameters. The design rationale is as follows: 1) By introducing the non-linear hyperbolic tangent function to adjust $K_{p}$, the Lagrange multiplier can respond quickly to large errors while avoiding frequent adjustments and reducing oscillations when the cost is close to the limit. When the average cost exceeds the limit ($\bar{c}>c_{th}$), $K_{p}$ and $K_{i}$ are increased to enhance the sensitivity of the PID controller to accelerate error correction. On the other hand, when the average cost is below the limit ($\bar{c}<c_{th}$), $K_{p}$ and $K_{i}$ are decreased to prevent overreaction of the PID controller, thus avoiding unnecessary oscillations and ensuring stability. By dynamically adjusting $K_{p}$ and $K_{i}$ in this manner, the controller adapts its low-frequency gain to match the current error magnitude, balancing speed and stability in error correction. 2) Because $K_{d}$ captures the volatility of cost changes, it will be adjusted using the standard deviation $\sigma_{c}$ of the cost over the period to stabilize cost control. A larger $\sigma_{c}$ indicates greater fluctuations in the error signal, suggesting the presence of high-frequency disturbances or noise. $K_{d}$ will be correspondingly increased, such that the controller can enhance its damping characteristics, adding phase lead and improving transient response. This helps mitigate the effects of sudden changes and stabilize the system.
@@
In a nutshell, combining VPA and aPID leads to our proposed framework Marvel: 1) Given the pretrained policy and Q-functions from offline learning, Marvel first applies VPA to align the pretrained Q-functions for both reward and cost using the offline dataset; 2) Marvel next utilizes Lagrangian-based online safe RL algorithms to further finetune both the pretrained policy and aligned Q-functions, by using aPID to update the Lagrange multipliers. We present the algorithmic framework of Marvel in Appendix A. Here VPA and aPID work in concert to guarantee the superior performance of Marvel: aPID addresses the Lagrange multiplier mismatch problem and quickly pushes down the potential high cost resulted by VPA, whereas VPA facilitates active exploration of high-reward state-action pairs and the usage of the pretrained policy as a warm-start for fast online finetuning with aPID control.

### v3

@@ 3 O2O Safe RL with Value Pre-Alignment
As shown in Fig. 1, naively finetuning the offline policy for safe RL would not work and the finetuned policy shows clear “inertia" in improving its performance: within a long period after online finetuning starts, its cost stays far below the limit, but its reward is quite low and not improving. This implies that such a strategy automatically “inherits" the conservatism from offline safe RL and is reluctant to actively explore to fully utilize the safe gap below the cost limit. In this section, we delve into this failure of naive finetuning, which points to two unique challenges for policy finetuning in O2O safe RL, i.e., erroneous offline Q-estimations and Lagrange multiplier mismatch. To address these, we propose a framework for O2O safe RL, i.e., warM-stArt safe RL with Value prE-aLignment (Marvel).
@@ 3.1 Pre-Finetune Phase
Challenge I: Erroneous Q-estimations resulting from objective mismatch and offline cost sparsity.  By learning from a fixed dataset only, offline safe RL typically suffers from large extrapolation errors for OOD actions beyond the support of the dataset. A general principle to handle this is to penalize the reward/cost estimations for the OOD actions in such a way that risky explorations outside the dataset are discouraged. The Q-functions can be optimized as follows:
$Q\text{:}\$$\min\mathbb{E}_{{D}}\Big[\big(Q(s,a)-(r+\gamma\mathbb{E}_{a^{\prime}\sim\pi_{\text{eval}}(\cdot|s^{\prime})}[Q(s^{\prime},a^{\prime})])\big)^{2}\Big]+\psi\cdot\mathcal{P}(s,a_{OOD}),$ (6)
$Q_{c}\text{:}\$$\min\mathbb{E}_{{D}}\Big[\big(Q_{c}(s,a)-(c+\gamma\mathbb{E}_{a^{\prime}\sim\pi_{\text{eval}}(\cdot|s^{\prime})}[Q_{c}(s^{\prime},a^{\prime})])\big)^{2}\Big]-\psi_{c}\cdot\mathcal{P}_{c}(s,a_{OOD}).$ (7)
@@
Here $\pi_{\text{eval}}$ denotes policy under evaluation. $\psi\cdot\mathcal{P}(s,a_{OOD})$ and $\psi_{c}\cdot\mathcal{P}_{c}(s,a_{OOD})$ are the penalty terms. For instance, penalties are introduced in VOCE (Guan et al., 2024) to minimize the reward Q-values and maximize the cost Q-values for OOD actions. CPQ (Xu et al., 2022) increases the perceived cost of OOD actions during Q-function and policy updates. In contrast, the optimization of Q-functions in online safe RL is standard without any penalty terms, i.e, $\psi=0$ in Eq. 6 and Eq. 7.
@@
Obviously, Offline and online safe RL have distinct objectives for Q-functions, meaning pretrained Q-functions may not accurately estimate values for state-action pairs encountered during online interactions. As a result, offline policies tend to act overly conservatively, exploring only low-cost regions during online finetuning. This limitation actually arises from the joint effect of (i) sparse offline costs—leading the pretrained cost critic $Q_{c}$ to assign low costs to IND actions—and (ii) offline pessimism that assigns high costs to OOD actions through conservative extrapolation. Under a Lagrangian update, the policy thus prefers to stay within the low-cost IND region and avoids OOD regions that the critic deems high-cost, even when those regions contain high rewards. However, effective online learning requires identifying state–action pairs with both high rewards and low costs, which necessitates exploring areas with potentially higher costs. This objective mismatch is even more pronounced in O2O safe RL due to the additional cost Q-function.
@@
Solution: Value Pre-Alignment.  To address the first challenge, a naive approach is to reevaluate the offline policy $\pi_{0}$ in online environments using Monte Carlo simulations, which however introduces additional interaction costs. Motivated by the recent advances in Off-Policy Evaluation (OPE) (Uehara et al., 2022), we borrow the idea from Fitted Q Evaluation (Hao et al., 2021) to align the offline Q-functions with the online learning objectives for the offline policy, by using the offline dataset before online policy finetuning. Starting from the pretrained Q-functions, we seek to minimize:
@@
$\mathcal{L}_{Q}^{\mathrm{VPA}}(\omega)$ $=\mathbb{E}_{D}\!\left[\!\Big(Q(s,a)-(\gamma\,\mathbb{E}_{a^{\prime}\sim\pi_{0}(\cdot|s^{\prime})}[\hat{Q}(s^{\prime},a^{\prime})-\alpha^{\mathrm{VPA}}\log\pi_{0}(a^{\prime}|s^{\prime})]+r)\Big)^{2}\right]$ (8)
$\mathcal{L}_{Q_{c}}^{\mathrm{VPA}}(\omega_{c})$ $=\mathbb{E}_{D}\!\left[\!\Big(Q_{c}(s,a)-(\gamma\,\mathbb{E}_{a^{\prime}\sim\pi_{0}(\cdot|s^{\prime})}[\hat{Q}_{c}(s^{\prime},a^{\prime})-\alpha_{c}^{\mathrm{VPA}}\log\pi_{0}(a^{\prime}|s^{\prime})]+c)\Big)^{2}\right]$ (9)
where $\omega$ denotes parameters of Q function and $\hat{Q}$ is target Q function. Here the entropy terms can result in both higher rewards and costs for state-action pairs with high entropy, where pretrained policy is ‘uncertain’. Offline policy remains unchanged during VPA to preserve knowledge extracted from offline data.
Assume that the mismatch between the offline policy-induced distributions of the learned policy $\pi_{0}$ and the offline behavior policy $\mu$ is bounded by constant $C<\infty$, i.e., $\max_{(s,a)\in S\times A}\frac{d^{\pi_{0}}\left(s,a\right)}{d^{\mu}\left(s,a\right)}\leq C$. Let $Q_{0}$, $Q_{K}$, and $\hat{Q}^{\pi_{0}}$ be Q-function obtained from offline phase, Q-function after $K$ iterations in VPA and the optimal (fixed-point solution) Q-function for reward of $\pi_{0}$, respectively (similarly, $Q_{c,0}$, $Q_{c,K}$, and $\hat{Q}_{c}^{\pi_{0}}$ for cost Q-functions). We can have following result to justify effectiveness of VPA:
Theorem 1.
Assume that the reward $r$ and cost $c$ are bounded by $1$ for all $(s,a)\in S\times A$. With probability at least $1-\delta$, the Q-estimation errors in VPA can be bounded in the weighted $L_{2}$-norm $\|\cdot\|_{2,d^{\pi_{0}}}$ under the state-action distribution induced by policy $\pi_{0}$:
$\|Q_{K}-\hat{Q}^{\pi_{0}}\|_{2,d^{\pi_{0}}}\leq\frac{\sqrt{C\tilde{\epsilon}}}{1-\gamma}+\gamma^{K}\|Q_{0}-\hat{Q}^{\pi_{0}}\|_{2,d^{\pi_{0}}},$~ (10)
$\|Q_{c,K}-\hat{Q}_{c}^{\pi_{0}}\|_{2,d^{\pi_{0}}}\leq\frac{\sqrt{C\tilde{\epsilon}_{c}}}{1-\gamma}+\gamma^{K}\|Q_{c,0}-\hat{Q}^{\pi_{0}}_{c}\|_{2,d^{\pi_{0}}}$ (11)
where $\tilde{\epsilon}$ and $\tilde{\epsilon}_{c}$ are the approximation errors. It decreases as the size of the dataset increases while increases with the values of $\alpha^{VPA}$ and $\alpha_{c}^{VPA}$, respectively.
The full proof and the specific expression of $\tilde{\epsilon}$ are provided in Appendix H. As shown in Theorem 1, the overall estimation error consists of two terms, i.e., the statistical error in the first term and the iteration error in the second term. With a larger dataset and more expressive function approximator, the first term vanishes, and the overall error is dominated by the iteration error, which decays exponentially with training steps $K$. This indicates that VPA can yield increasingly accurate Q-function estimates. Moreover, since reward optimization dominates policy updates at the early stage of online finetuning when Lagrange multiplier is small, a larger uncertainty in reward Q-estimations is preferred to facilitate more online exploration. To this end, Theorem 1 further implies that a larger $\alpha^{VPA}$ should be selected in Eq. 8 for reward Q-values, whereas $\alpha_{c}^{VPA}$ in Eq. 9 should be smaller to ensure a more accurate estimation of cost Q-values. While the theory suggests improved Q estimations for IND and OOD actions, our empirical findings show estimations for OOD actions after VPA tends to be overoptimistic. This discrepancy is likely due to imperfect data coverage and function approximation errors. Interestingly, while such overestimation is problematic and avoided for offline RL without online explorations, it is indeed beneficial for O2O RL during early phase of online finetuning, which can encourage more active explorations beyond the offline dataset. While Luo et al. (2024) considered a similar analysis for unconstrained RL, we focus on the safe RL setting and further introduce VPA, a pre-finetuning procedure incorporating an entropy term to mitigate erroneous Q-value estimations.
@@
Table 1: Spearman’s rank correlation coefficients of $Q$-value and $Q_{c}$-value in BallCircle and CarRun. “Random” refers to rollouts starting from randomly initialized state-action pairs (potentially OOD), while “Dataset” refers to rollouts from state-action pairs sampled from the offline dataset.
@@
To empirically characterize the performance of VPA in correcting Q-estimations, we leverage Spearman’s rank correlation coefficient, which measures the strength and direction of a monotonic relationship between two ranked variables. The reason is that relative ranking of Q-values are more important than absolute values for policy update. We provide more details of Spearman’s rank correlation coefficient in Appendix G.1. Specifically, given a dataset collected by the offline policy in the environment, we compare the ranking of learned reward/cost Q-values before and after VPA with that of estimated actual return via Monte Carlo simulations. A large Spearman’s rank correlation coefficient implies that distribution of learned Q-values is more aligned with the distribution of true Q-values. As shown in Table 1, it is evident that coefficient increases significantly after VPA for both reward and cost Q-values, no matter if offline policy rolls out from a seen state-action pair in offline dataset or from a randomly selected OOD state-action pair. This clearly demonstrates effectiveness of VPA in aligning the pretrained Q-functions.
@@ 3.2 Finetune Phase
Challenge II: Lagrange multiplier mismatch.  Conventional value-based online safe RL relies on updating Lagrange multipliers alongside the policy and Q-functions during training, to push the overall cost below the limit while balancing between maximizing the reward and minimizing constraint violations. Although the policy and Q-functions can benefit from offline pretraining for a warm start, offline safe RL algorithms like CPQ (Xu et al., 2022) and BEAR-lag (Ray et al., 2019) cannot accurately estimate Lagrange multipliers with regularizing strengths matching the cost of the offline policy, e.g., a small multiplier is not powerful enough to push down the policy cost, while a large multiplier prevents active exploration of high-reward state-action pairs. For instance, in BallCircle, the offline Lagrange multiplier value obtained using the BEAR-lag algorithm is approximately 1500, whereas during online finetuning, the SAC-lag requires a value of only about 0.65. The gap between these values precludes the direct use of offline Lagrange multipliers. Improper initialization can lead to extensive constraint violations or training stagnation, an issue we term as the Lagrange multiplier mismatch.
@@
On the other hand, as VPA promotes active exploration of high-reward state-actions by optimistically estimating rewards and pessimistically estimating costs, it inevitably increases the risk of exploring high cost state-actions, which in turn amplifies the need for appropriate Lagrange multipliers to quickly reduce the constraint violations.
Figure 2 compares online finetuning performance in BallCircle when using VPA with: (1) a well-initialized Lagrange multiplier vs. (2) zero initialization, both with standard dual ascent updates. A good initial multiplier enables effective cost control, while poor initialization leads to significant constraint violations and slower convergence. The results also imply that although VPA aligns the distributions of Q-values, it may introduce high costs for online finetuning, which can be addressed with an appropriate initial Lagrange multiplier.
Solution: Adaptive PID Control.  Finding a good initial value of the Lagrange multiplier can address the mismatch problem and mitigate the risk brought by VPA. However, obtaining such a value by manual tuning is challenging, and no theory provides a reliable prediction. We therefore adapt the multiplier rapidly and effectively by using a PID-style controller that has shown strong empirical performance in online safe RL (Stooke et al., 2020; Yuan et al., 2022; Zhou et al., 2021), and we tailor it to the O2O setting.
We use a discrete, position-form PID update. Let the discounted total episode cost be $c_{t}$ at training iteration $t$, the limit be $c_{th}$, and the violation be $e_{t}=c_{t}-c_{th}$. Define an EMA-smoothed violation $\tilde{e}_{t}=a_{p}\tilde{e}_{t-1}+(1-a_{p})e_{t}$ with smoothing factor $a_{p}\in[0,1)$. Define an EMA-smoothed cost $\tilde{c}_{t}=a_{d}\tilde{c}_{t-1}+(1-a_{d})c_{t}$ with smoothing factor $a_{d}\in[0,1)$, and a derivative buffer of length $d\in\mathbb{N}$. The integral state and the derivative signal follow the implementation:
$I_{t}\;=\;\max\!\big(0,\;I_{t-1}+K_{i}\,e_{t}\big),\qquad D_{t}\;=\;\max\!\big(0,\;\tilde{c}_{t}-\tilde{c}_{t-d}\big).$
The multiplier is then
$\lambda_{t}\;=\;\max\!\big(0,\;K_{p}\,\tilde{e}_{t}+I_{t}+K_{d}\,D_{t}\big).$ (12)
This position-form matches the code path where the proportional term uses $\tilde{e}_{t}$, the integral term uses $e_{t}$, and the derivative term uses a lagged difference of smoothed costs $D_{t}$. The integral channel provides the same role as dual ascent.
VPA relaxes offline conservatism and increases exploration of high-return state–action pairs during early finetuning, which can raise violations. A larger control strength in early stages reduces violations rapidly, and a smaller control strength near the limit stabilizes learning and suppresses oscillations. This motivates an adaptive gain schedule.
Gains are adapted from a short buffer of smoothed costs. Let the buffer contain the last $n$ values of $\tilde{c}_{t}$ that are also used to form the derivative buffer. Define the average $\bar{c}_{t}=\frac{1}{n}\sum_{i=1}^{n}\tilde{c}_{t+1-i}$ and the standard deviation $\sigma_{t}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\tilde{c}_{t+1-i}-\bar{c}_{t})^{2}}$. With coefficients $\alpha,\beta,\gamma>0$, the update is
@@
$K_{p}$ $\leftarrow\mathrm{clip}\!\left(K_{p}\cdot\left(1+\alpha\cdot\frac{\bar{c}_{t}-c_{th}}{\bar{c}_{t}}\right),\;K_{p,\min},\;K_{p,\max}\right),$ (13)
@@
$K_{i}$ $\leftarrow\mathrm{clip}\!\left(K_{i}\cdot\left(1+\beta\cdot\frac{\bar{c}_{t}-c_{th}}{\bar{c}_{t}}\right),\;K_{i,\min},\;K_{i,\max}\right),$ (14)
$K_{d}$ $\leftarrow\mathrm{clip}\!\left(K_{d}\cdot\left(1+\gamma\cdot\frac{\sigma_{t}}{\bar{c}_{t}}\right),\;K_{d,\min},\;K_{d,\max}\right).$ (15)
This schedule increases $(K_{p},K_{i})$ when $\bar{c}_{t}>c_{th}$ to raise low-frequency gain and accelerate correction, and decreases $(K_{p},K_{i})$ when $\bar{c}_{t}<c_{th}$ to avoid overreaction and improve stability. The schedule increases $K_{d}$ when $\sigma_{t}$ is large to provide additional damping against short-horizon volatility. $K_{d}$ tends to plateau naturally once $\sigma_{t}$ stabilizes. The minimum and maximum bounds of each PID gain are set to 0.1 and 10 times its initial value, respectively. Since $K_{d}$ is non-decreasing by design, we set $K_{d,\min}$ only to keep notation consistent with $(K_{p,\min},K_{i,\min})$.
Appendix D.3.2 demonstrates that these parameters are easy to tune but also robust in their own selection (e.g., same across all environments in the experiments).
@@
In a nutshell, combining VPA and aPID leads to our proposed framework Marvel: 1) Given the pretrained policy and Q-functions from offline learning, Marvel first applies VPA to align the pretrained Q-functions for both reward and cost using the offline dataset; 2) Marvel next utilizes Lagrangian-based online safe RL algorithms to further finetune both the pretrained policy and aligned Q-functions, by using aPID to update the Lagrange multipliers. We present the algorithmic framework of Marvel in Appendix A.
Here VPA and aPID work in concert: aPID addresses the Lagrange multiplier mismatch problem and quickly pushes down the potential high cost resulted by VPA, whereas VPA facilitates active exploration of high-reward state-action pairs and the usage of the pretrained policy as a warm-start for fast online finetuning with aPID control.